OpenStudy (anonymous):

Charles's Law of V/T = k were represented in graphical form, what it would look like?j

OpenStudy (unklerhaukus):

do you want to plot \(V(T)=kT\) or \(T(V)=\frac Vk\)

OpenStudy (anonymous):

t(v) = v/k.

OpenStudy (unklerhaukus):

is of the form\[y(x)=mx+b\] where \((x,y)=(V,T)\) the vertical-intercept is zero the gradient is \(1/k\)